# Do We Need Subsidiarity in Software?

### v1

@@ 6.3. What other attitudes emerged from the interviews?
Desiring platform control under better conditions embodies the sentiment that our participants would like the benefits of having platforms control their data in an ideal world, but do not trust “big tech” to do so responsibly. Some participants expressed a desire for greater regulation on how platforms use their data such that they could trust a platform level of control. P13 described it this way: “I might think differently if there were stronger safeguards, where people didn’t have to worry…I think that government regulation should be the goal.” All participants who expressed this sentiment ultimately chose a user level of control. Desiring platform control under better conditions was mentioned by 2 participants for operating systems, 2 participants for browsers, 1 participant for email, 3 participants for search engines, 1 participant for social media, 1 participant for chat applications, and was mentioned by 8 unique participants.
User-level preference but platform-level necessity embodies an opposite sentiment to the previous theme: in this case, participants desired the user level of control of their data but thought the platform level was necessary for its benefits, such as making platforms free. P2 expressed this sentiment as, “ideally [control] would be at the user level, but I just don’t think that’s realistic in the current [economic] model.” User-level preference but platform-level necessity was mentioned by 2 participants for browsers, 2 participants for email, 2 participants for search engines, 2 participants for social media, 1 participant for chat applications, and was mentioned by 2 unique participants.
Limiting data controlled at the platform level is the sentiment that while it’s important to have some data be controlled at the platform level, this data should be limited in scope, for example data about operating system crashes, or should require informed consent before sharing. P12 said, “some of the data I don’t mind [sharing]…the system operations, technical stuff.” Limiting data controlled at the platform level was mentioned by 4 participants for operating systems, 2 participants for browsers, 2 participants for search engines, 2 participants for social media, and was mentioned by 6 unique participants.
Limitations of community control encompasses participant concerns with the community level of control even when selecting that as the necessary level. Some participants particularly highlighted concerns with worker surveillance when the community in question is an employer. Limitations of community control was mentioned by 2 participants for operating systems, 1 participant for browsers, 1 participant for search engines, 1 participant for social media, 1 participant for chat applications, and was mentioned by 3 unique participants.
@@ 7. Comparison of Actual Levels of Control and User-Identified Necessary Levels of Control
In this section, we will discuss if subsidiarity is upheld for the different software application types we studied based on the levels of control we determined in the data flow monitoring and what level of control our interview participants thought was the necessary level for that type of software. Although our data flow monitoring was only able to distinguish “lower” and “higher” levels of control, for the purpose of comparison we will equate lower levels of control determined from data flow monitoring with user-level control, and higher levels of control determined from data flow monitoring with platform-level control. We will only consider interview participant responses for necessary level of control in the personal usage case, since our data flow monitoring did not detect data flows to community organizations. (We will address participant desires for community-level control in the discussion.) Although a majority of our interview participants wanted user-level control for all software types, we distinguish between cases of a strong majority (12-16 participants wanting user-level control) and a weak majority (8-11 wanting user-level control, and 4-7 wanting platform-level control). In the cases of weak majority, because more participants wanted platform level of control compared to the strong majority, we discuss options for both user and platform level of control.
@@ 8. Discussion
We discuss the design implications of our findings, including how the principle of subsidiarity can inform what to build, the implications of greater trust in communities over corporations, limitations, and opportunities for future work.
@@ 8.1. Design Implications: Violations of Subsidiarity Can Show Us What To Build
The principle of subsidiarity helps us discern when control should be in the hands of the user and when it is necessary to move control to a higher level to achieve some better outcome for users. The level of control that a majority of users identify as necessary for a particular type of software can inform what types of products to build: both when to prioritize giving users full control over their data, and when users may be okay with automatically sharing their data in return for better features. In light of this, we will discuss two software types with particularly salient findings from this study: chat applications and search engines.
@@ 8.1.1. Improving Chat Applications
Chat applications were the only software type where our participants unanimously wanted user-level control for personal usage. Yet, neither of the chat applications we considered offered user-level control, and Slack was the most egregious across all software types in contacting advertising platforms. Furthermore, one of the themes for why participants wanted privacy in chat applications (as well as email and social media) was because of an expectation of privacy because of the design and purpose of the application. In other words, because chat applications make the user think they are having a private conversation, there is a greater expectation of privacy in comparison to applications like search engines. While some chat platforms not included in our study may likely offer a user-level of control (for example Telegram, especially with its strictest privacy controls), our interview results indicate that it is unacceptable that any chat application would not have strong protections over user privacy. Chat applications should be better designed to protect user privacy by default, following design guidelines such as Privacy by Design (Cavoukian et al., 2009), or at the very least require informed consent or have greater transparency if user data will be collected. An element of this transparency — for all platforms, not just chat applications — is related to the platform’s design. Platforms designed in ways that make users think they are private should not collect data unexpectedly, and conversely, platforms that will collect data should be designed in ways that give the user this expectation.
@@ 8.1.2. Search
With search engines, we found a situation almost opposite to that with chats. In the case of search engines, more of our participants were open to a platform level of control. However, two out of the three search engines that we evaluated operated at a user level of control. While it is still important for these platforms to exist to cater to those who want a user-level of control, there is a discrepancy: more users find a platform level of control necessary for search engines compared to chat applications, yet there are multiple options for search engines that offer a user level of control. Considering which software types have the most votes for the user-level of control can help us prioritize which software types to build more privacy-focused alternatives for.
@@ 8.2. Trust in Communities Over Corporations
Our interviews revealed a greater trust in community organizations — whether these are workplaces, schools, or special interest groups — over “big tech” corporations. In many cases for our participants, the question was not if they wanted their data controlled at a level higher than the user level, the question was who is this higher level? When discussing community contexts, our participants highlighted that the community providing the software to them has a right to control their data. However, they predominantly did not mention such a right when discussing the technology company that built the software, even though the software is, in a sense, being provided to them by the company. This underscores a difference in understanding of the rights of those developing the technology and those administrating the technology in terms of who has access to data. In addition to a different understanding of rights of access, there were also differing levels of trust in community organizations compared to tech companies. Our participants frequently mentioned a distrust of large tech companies, but also frequently mentioned they trust community-level institutions because of their different incentives and because they operate at a smaller scale. Additionally, a few participants mentioned a desire for something that we don’t currently see existing: community collectives to help protect security or provide higher-level features typically provided by tech companies and powered by automatic data collection. Our participants want these higher-level features but are more comfortable sharing their data at the community level to receive them. We encourage this as a new direction.
8.3. Limitations and Future Work
@@
Data flow monitoring comes with a number of inherent limitations. Because the packets were encrypted, we were not able to read their contents. Thus, when we observed data transfers, we could not say for sure if the data is sensitive user data or not. We did our best to mitigate this by only considering larger packets, but the approach is not foolproof. Nevertheless, we hope that even if our data analysis approach is not absolutely precise, it still can provide a general sense of what is going on.
@@
Another limitation lies in the software we studied, especially in certain categories. For example, in the chat application category we did not observe any applications that operated at a lower level of control, but our results may have been different if we included additional applications, such as Signal or Telegram. We recommend future work investigating more software platforms, especially chat applications.
@@
Finally, our interview sample consisted of graduate students and employees from our university. Thus, everyone in our participant group holds at least a bachelors degree, and many of our participants hold masters degrees or higher. While a higher level of education among our participants may have increased participants’ familiarity with privacy concerns in technology and thus likely allowed them to give more informed responses, it also may have biased their responses. Additionally, recruiting from our university meant that all interview participants currently reside in the Midwestern United States. While many of our participants were not originally from the Midwest, and some were not originally from the United States, there are still potential perspectives lost from this sample that would have surfaced in a more globally diverse interview pool. Given that perspectives of privacy can differ cross-culturally (Trepte et al., 2017), there is value in future work that investigates if cultural differences impact an understanding of which level of control is necessary in different technologies.
@@ 9. Conclusion
Appendix A A Deeper Dive Into Unexpected Endpoints Contacted During Activity
To determine the level of control an application operated at, we only considered if application data was transferred while the software ran idle, and if any unexpected endpoints were contacted at all. Thus, the application data transferred to unexpected endpoints was not a determining factor in the level of control. Yet, it is still interesting to consider how much application data was transferred to unexpected endpoints, to determine if they are collecting user data.
In our analysis approach in the paper, we focused on only transfers of application data over 1000 bytes, as transfers of application data under 1000 bytes were likely to be encrypted keep-alives or heartbeats. However in the case of unexpected endpoints, we observed that while most of the application data transfers in these cases were under 1000 bytes, many had at least one data transfer between 350 and 1000 bytes, and very few of the data transfers resembled the regular patterns of application data transfer observed for heartbeats or keep alives (e.g. every 30 or 60 seconds). Thus, although the application data transfers in these cases were under our threshold, we report them as they may not be simple keep alives. For these cases, we report the number of data transfers to that unexpected IP and the largest application data transfer size in number of bytes. We report these results in Figure 17.
Figure 17. The applications that had data transfers to unexpected endpoints while taking an action, with the number of transfers of application data to that endpoint and the size in bytes of the largest application data transfer to that endpoint.
[…]
A table with five columns: Application, Browser, Organization, Number of Transfers of Application Data, and Largest Application Data Transfer Size (in bytes). The table rows detail data transfers for various applications: Gmail on Chrome, with Amazon as the organization. There were 4 transfers, with the largest size being 355 bytes. Instagram is shown with transfers on multiple browsers and organizations. On Chrome, transfers occurred with GetLinks SMC (5 transfers, 884 bytes), Horizon Scope Mobile (1 transfer, 130 bytes), and Reliance Jio (0 transfers, 0 bytes). On Firefox, transfers occurred with Coscom Reliability (7 transfers, 785 bytes). On Brave, transfers occurred with Telecom Algeria (4 transfers, 481 bytes). TikTok on Brave, with Google as the organization. There were 9 transfers, with the largest size being 399 bytes. X on Brave, with Google as the organization. There were 0 transfers, with the largest size being 0 bytes. Slack is shown with transfers on both Chrome and Firefox. On Chrome, transfers occurred with Microsoft (24 transfers, 568 bytes), Rubicon Project (2 transfers, 731 bytes), AS Criteo (4 transfers, 594 bytes), Rocketfuel (2 transfers, 1038 bytes), Rhythmone (4 transfers, 612 bytes), Appnexus (4 transfers, 583 bytes), and Facebook (0 transfers, 0 bytes). On Firefox, transfers occurred with Microsoft (9 transfers, 405 bytes), Rubicon Project (3 transfers, 609 bytes), AS Criteo (6 transfers, 479 bytes), Rocketfuel (4 transfers, 876 bytes), AS Outbrain (7 transfers, 672 bytes), and Appnexus (5 transfers, 439 bytes). Discord is shown with transfers on multiple browsers. On Firefox, transfers occurred with Microsoft (5 transfers, 158 bytes). On Brave, transfers occurred with Google (0 transfers, 0 bytes).
We observe from Figure 17 that Slack transfers application data to the most unexpected endpoints (6 per browser experiment), and the majority of the largest application data transfers from Slack are between 400 and 1050 bytes. Instagram transfers application data to unexpected foreign (from the perspective of the United States) endpoints, and especially has some larger (700-900 bytes) application data transfers to GetLinks SMC (based in Pakistan) and Coscom Reliablity (based in Uzbekistan). On the other end of the spectrum, X and Discord both contacted Google on Brave but transferred 0 bytes of application data.
Although all but one of the largest application data transfer were under our threshold of 1000 bytes, because many of the largest transfers were between 500 and 1000 bytes on both Instagram and Slack, and because Instagram and Slack had the most unexpected endpoints, we highlight these two platforms as platforms that particularly exercise a higher level of control.

### v2

@@ 6.3. What other attitudes emerged from the interviews?
Desiring platform control under better conditions embodies the sentiment that our participants would like the benefits of having platforms control their data in an ideal world, but do not trust “big tech” to do so responsibly. Some participants expressed a desire for greater regulation on how platforms use their data such that they could trust a platform level of control. P13 described it this way: “I might think differently if there were stronger safeguards, where people didn’t have to worry…I think that government regulation should be the goal.” All participants who expressed this sentiment ultimately chose a user level of control.
User-level preference but platform-level necessity embodies an opposite sentiment to the previous theme: in this case, participants desired the user level of control of their data but thought the platform level was necessary for its benefits, such as making platforms free. P2 expressed this sentiment as, “ideally [control] would be at the user level, but I just don’t think that’s realistic in the current [economic] model.”
Limiting data controlled at the platform level is the sentiment that while it’s important to have some data be controlled at the platform level, this data should be limited in scope, for example data about operating system crashes, or should require informed consent before sharing. P12 said, “some of the data I don’t mind [sharing]…the system operations, technical stuff.”
Limitations of community control encompasses participant concerns with the community level of control even when selecting that as the necessary level. Some participants particularly highlighted concerns with worker surveillance when the community in question is an employer. P13 said, “I would be skeptical of an employer wanting to monitor by the minute employee productivity so that they can figure out ways to extract more labor that are potentially dehumanizing.”
@@ 7. Comparison of Actual Levels of Control and User-Identified Necessary Levels of Control
In this section, we will discuss if subsidiarity is upheld for the different software application types we studied, comparing our data flow monitoring results to the necessary level of control identified in the interviews. Although our data flow monitoring was only able to distinguish “lower” and “higher” levels of control, for comparison we will equate “lower” control with user-level control, and “higher” control with platform-level control. We only consider interview participant responses for necessary level of control in the personal usage case, since we did not detect data flows to community organizations. (We will address community-level control in the discussion.) Although a majority of our interview participants thought user-level control was necessary for all software types, we distinguish between cases of a strong majority (12-16 participants selecting user-level control) and a weak majority (8-11 selecting user-level control). In the cases of weak majority, we discuss options for both user and platform level of control.
@@ 8. Discussion
We discuss the design implications of our findings, including how the principle of subsidiarity can inform what to build, the technical and logistical challenges of implementing user-level control, the implications of greater trust in communities over corporations, limitations, and opportunities for future work.
@@ 8.1. Design Implications: Violations of Subsidiarity Can Show Us What To Build
Given that willingness to share data can be contextual (Hudig and Singh, 2025) and that some users find data-sharing to be useful (Ur et al., 2012), the principle of subsidiarity helps us discern when control should be in the hands of the user and when it is necessary to move control to a higher level to achieve some better outcome for users. In light of this, we will discuss two software types with particularly salient findings from this study: chat applications and search engines.
@@ 8.1.1. Improving Chat Applications
Chat applications were the only software type where our participants unanimously found user-level control necessary for personal usage. Yet, neither of the chat applications we considered offered user-level control, and Slack was the most egregious across all software types in contacting advertisers. Furthermore, one of the themes for why participants found privacy necessary in chat applications was because of an expectation of privacy because of the design and purpose of the application. In other words, because chat applications make the user think they are having a private conversation, there is a greater expectation of privacy in comparison to applications like search engines. While some chat platforms not included in our study may offer user-level control (for example Telegram, especially with its strictest privacy controls), our interview results indicate that it is unacceptable that any chat application would not have strong protections over user privacy. Especially given that a majority of users have negative feelings towards data tracking but don’t take any steps to minimize tracking themselves (Coopamootoo et al., 2022), chat applications should be better designed to protect user privacy by default. This could be achieved by following design guidelines such as Privacy by Design (Cavoukian and others, 2009) or requiring informed consent. In light of prior findings that users make different choices with their data when offered greater transparency (Liccardi et al., 2014; Van Kleek et al., 2017), a platform’s interface design can help achieve such transparency. Platforms designed in ways that make users think they are private (such as chat interfaces) should not collect data unexpectedly, and conversely, platforms that will collect data should be designed in ways that give the user this expectation.
@@ 8.1.2. Search
With search engines, we found a situation almost opposite to that with chats. In the case of search engines, more of our participants were open to platform-level control. However, two out of the three search engines that we evaluated operated at user-level control. While it is still important for these platforms to exist to cater to those who want user-level control, there is a discrepancy: more users find platform-level control necessary for search engines compared to chat applications, yet there are multiple options for search engines that offer user-level control. Considering which software types have the most votes for user-level control can help us prioritize which software types to build more privacy-focused alternatives for.
8.2. Technical and Logistical Challenges of Implementing User-Level Control
Even with available options at desired levels of control, technical or logistical challenges may exist with employment, particularly at the user level. For example, someone may desire an operating system with user-level control but not have the technical ability to install an OS like Linux. Similarly, someone may prefer an email service with user-level control but be forced to use a platform like Gmail by their employer. Thus, we cannot achieve subsidiarity only by having a myriad of options to choose from: other systemic changes are required. An example of this direction is software like Linux Mint (32), a Linux distribution that works “Out of the Box” but is still highly configurable, providing user-level control in an accessible way.
We focused on subsidiarity at the data collection stage: in the current technical landscape it is harder to control what a platform does with data once they have it, so collection is the first level of intervention. However, a more robust notion of subsidiarity would include the usage and retention stages of the data life cycle. User-level control of usage might look like Blockchain technologies that allow stakeholders to have finer-grained control over their data after it is collected (Mann et al., 2021). User-level control of retention would entail mechanisms that ensure data deletion (Ng, 2019).
An emerging challenge to user-level control is the integration of LLM-based assistants in everyday applications, for example, the “AI Overview” in Google Search. There are pervasive public anxieties that LLMs are being trained on email and word processor data (Westerholm, 2025; McCurdy, 2025). While tech companies deny that they actually do this (Westerholm, 2025; McCurdy, 2025), these concerns highlight that more data is needed to train LLMs than is required for targeted advertising or general user experience improvements, thus exacerbating already-existing concerns with data collection. Although data collection for LLMs was an example mentioned by the interviewer in our thought exercises, our participants spoke more about targeted advertising. Yet, LLMs have relevance to many of the themes justifying user- or platform-level control; for example, platform-level control for improved user experience or user-level control for increased privacy. While our findings generally apply to the case of automatic data collection for LLM training, user-level control may be trickier to ensure in conversational interfaces: users may disclose more information to LLMs than on other platforms due to LLMs’ conversational interface and imitation of human empathy (Ali et al., 2025). While disclosing personal information to a conversational LLM technically falls under user-level control because the user is choosing to give up this data, is it really user-level control if the interface is subtly coercing them to do so?
@@ 8.3. Trust in Communities Over Corporations
Prior work found that willingness to share data can be dependent on who it is shared with (Hudig and Singh, 2025). Building on this, our interviews revealed a greater trust in community organizations — whether these are workplaces, schools, or special interest groups — over “big tech” corporations. In many cases for our participants, the question was not if they wanted their data controlled at a level higher than the user level, the question was who is this higher level? When discussing community contexts, our participants highlighted that the community providing the software to them has a right to control their data. However, they predominantly did not mention such a right when discussing the technology company that built the software, even though the software is, in a sense, being provided to them by the company. This underscores a difference in understanding of the rights of those developing the technology and those administrating the technology in terms of who has access to data. In addition to a different understanding of rights of access, there were also differing levels of trust in community organizations compared to tech companies. Our participants frequently mentioned a distrust of large tech companies, but also frequently mentioned they trust community-level institutions because of their different incentives and because they operate at a smaller scale. Additionally, a few participants mentioned a desire for something that doesn’t currently exist: community collectives to help protect security or provide higher-level features typically provided by tech companies and powered by automatic data collection. Our participants want these higher-level features but are more comfortable sharing their data at the community level to receive them. We encourage this as a new direction.
8.4. Limitations of Study and Future Work
@@
Data flow monitoring comes with a number of inherent limitations. Because the packets were encrypted, we were not able to read their contents. Thus, when we observed data transfers, we could not say for sure if the data was sensitive user data or not. We did our best to mitigate this by only considering larger packets, but the approach is not foolproof. Nevertheless, we hope that even if our data analysis approach is not absolutely precise, it still can provide a general sense of what is going on.
@@
Another limitation lies in the software we studied, especially in certain categories. For example, in the chat application category we did not observe any applications that operated at a lower level of control, but our results may have been different if we included additional applications, such as Signal or Telegram. Additionally, although we only studied Windows and Linux because we could install these operating systems on laptops with the same hardware, Apple software products would provide an interesting comparison given their reputation of being more privacy-centric than other companies (Leswing, 2021). We recommend future work investigating more software platforms, especially chat applications and Apple products.
@@
Finally, our interview sample consisted of graduate students and employees from our university. Thus, everyone in our participant group holds at least a bachelors degree, and many of our participants hold masters degrees or higher. While a higher level of education among our participants may have increased participants’ familiarity with privacy concerns in technology and thus likely allowed them to give more informed responses, it also may have biased their responses. In particular, those with more privileged backgrounds do not suffer as many harms from data privacy violations as do members of marginalized groups (Sannon and Forte, 2022). Thus, there is an important opportunity for future work exploring how people who are marginalized perceive what the “necessary” level of control is, as this will likely differ in comparison to the population we interviewed. Additionally, recruiting from our university meant that all interview participants currently reside in the Midwestern United States. While many of our participants were not originally from the Midwest, and some were not originally from the United States, there are still potential perspectives lost from this sample that would have surfaced in a more globally diverse interview pool. Given that perspectives of privacy can differ cross-culturally (Trepte et al., 2017), there is value in future work that investigates if cultural differences impact an understanding of which level of control is necessary in different technologies.
@@ 9. Conclusion
Appendix A Interview Guide
The full interview protocol is as follows:
The interviewer first introduces the tradeoffs of giving control to higher or lower levels.
Claimed Benefits: Better understanding of user preferences to make a better product, better recommendation algorithms, allows for the product to be free, don’t have to make customization decisions (e.g. Apple’s design team may have better aesthetics than yours), able to share data with others for better cooperative work e.g. sharing calendar availability with coworkers, sharing location with friends and family.
Claimed Drawbacks: Loss of control over one’s digital identity, ability for sensitive data to end up in the wrong hands, corporate or government surveillance including exploitation of user data for targeted advertising, behavioral modification, less ability to customize the experience to your liking (window managers, custom status bars, “ricing”) and pushing from the company to use their products (e.g. Microsoft strongly pushing people to use the Microsoft edge browser, Microsoft and Apple forcing people to use their AI tools in apps).
Interviewer: Rate on a scale of 1 to 5, 1 being not important at all that I have full control over my data, 5 being really important that I have full control over my data, how much do you want to control your data on different types of platforms?
  • Your computer’s operating system
  • Your browser
  • Search engines
  • Email providers
  • Social media platforms
  • Chat platforms such as Slack or Discord
¡Note: the data collected from these questions was redundant with data collected later in the interview and thus is not reported in the paper¿
The interviewer then shares the findings from our data flow monitoring study with the participants.
Interviewer: Is any of this surprising to you? Are you concerned about any of this?
The interview then proceeds to the thought exercises and identifications of necessary levels of control.
Interviewer: Consider the operating system of your laptop.
Pretend you’re a software engineer at Microsoft working on the Windows OS team. Why might you think it’s necessary for Microsoft to control user data from the operating system? Examples shared if the participant struggled: crash reports/bug fixes, location data for weather services, personalized news offerings, collecting user data for product usage, collecting user data for AI training.
Interviewer: Pretend you are the organizer of a company, university, or special interest group that has provided laptops to all its members. Why might you think it’s necessary for you, the head of the organization, to control user data from the operating system?
Examples shared if the participant struggled: software license compliance/automatic software updates, measuring how the laptops are benefitting the members or which software applications are being used the most in order to understand how the members of the organization are using the laptops and if they need extra training for certain types of software.
Interviewer: Why might it be necessary for all the data from your operating system to be completely controlled by you, the user?
Examples shared if the participant struggled: privacy, especially because people likely keep a lot of sensitive information on their laptops.
Interviewer: Based on thinking through these different situations, what level do you think is simultaneously the lowest possible and highest level of control necessary for operating systems? Does this change in different contexts, e.g. personal or professional?
Interviewer: Consider a web browser like Google Chrome or Firefox. Pretend that you’re a software engineer at Google, working on Google Chrome. Why might you think it would be necessary for Google to collect and control user data? Examples shared if the participant struggled: advertising to make the platform free, improved user experience.
Interviewer: Pretend that you are the head of a company, university or special interest group that has provided laptops to all its members. Why might it be necessary for you, the head of the organization, to control user data from the browser?
Examples shared if the participant struggled: detecting phishing or guarding against unsafe downloads, compliance with group rules (e.g. no piracy or pornography), understanding how employees or members are using different web applications for targeted skill-building or digital literacy training
Interviewer: Why might it be necessary for a user to have full control over their browsing data? Examples shared if the participant struggled: data privacy – using the internet for personal reasons like looking up health information, advocating for oneself in a marginalized group, etc.
Interviewer: Based on thinking through these different situations, what level do you think is simultaneously the lowest possible and highest level of control necessary for web browsers? Does this change in different contexts, e.g. personal or professional?
Interviewer: Consider an email platform like Gmail. Pretend that you’re a software engineer at Google on the Gmail team. Why might you think that it would be necessary for Google to collect and control user data and use it for purposes other than sending email? Examples shared if the participant struggled: spam filtering, taking flight data from email and turning it into google calendar events, targeted advertising to keep the platform free.
Interviewer: Pretend that you are the head of a company, university or special interest group that has provided laptops to all its members. Why might it be necessary for you, the head of the organization, to control user email data?
Examples shared if the participant struggled: phishing and scam detection, monitoring engagement with group communications (e.g. seeing if members open and read emails that are sent out), automatic meeting scheduling based on email contents.
Interviewer: Why might it be necessary that the individual user have full control over their email account?
Examples shared if the participant struggled: data privacy, especially because emails can contain particularly sensitive data.
Interviewer: Based on thinking through these different situations, what level do you think is simultaneously the lowest possible and highest level of control necessary for email accounts? Does this change in different contexts, e.g. personal or professional?
Interviewer: Now consider a search engine like Google Search. Pretend that you’re a software engineer at Google. Why might you think that it would be necessary for Google to have control over user data?
Examples shared if the participant struggled: personalization for improved user experience, advertising that makes the platform free.
Interviewer: Pretend that you are the head of a company, university or special interest group that has provided laptops to all its members. Why might it be necessary for you, the head of the organization, to control user data from a search engine?
Examples shared if the participant struggled: blocking unsafe sites from search results, tailoring resources (if employees are searching the same problem there can be training sessions about it).
Interviewer: Why might it be necessary for the user to have full control over their search data?
Examples shared if the participant struggled: data privacy — people search for sensitive things.
Interviewer: Based on thinking through these different situations, what level do you think is simultaneously the lowest possible and highest level of control necessary for search engines? Does this change in different contexts, e.g. personal or professional?
Interviewer: Now consider a social media platform like Instagram, TikTok, YouTube, X, or Bluesky. Imagine you’re a software engineer for one of these platforms. Why might you think it is necessary to collect and control user data?
Examples shared if the participant struggled: targeted advertising that keeps the platforms free, making recommendation algorithms more targeted to the user.
Interviewer: Currently, social media platforms like Instagram or X operate at the company level of control. However imagine a social platform for, say, a special interest group or hobby community, that is controlled entirely by the leader of the group or another community admin. If you are the admin for this special interest group’s platform, why might it be necessary for you to control the group’s data?
Examples shared if the participant struggled: community moderation and safety, program engagement (see which hash tags, topics, or accounts users interact with the most to better tailor community discussions).
Interviewer: Why might it be necessary for a user to have full control over their social media data?
Examples shared if the participant struggled: data privacy.
Interviewer: Based on thinking through these different situations, what level do you think is simultaneously the lowest possible and highest level of control necessary for social media platforms? Does this change in different contexts, e.g. personal or professional?
Interviewer: Now consider a chat platform like Slack, Discord, Whatsapp, or iMessage. Imagine you’re a software engineer for one of these platforms. Why might you think it’s necessary to collect and control user data?
Examples shared if the participant struggled: selling data to advertisers to keep the platform free, crash reports.
Interviewer: Imagine you are the head of a company, university, or special interest group that uses a chat platform like Slack or Discord for its internal communications. Why might it be necessary for you to control the group’s data?
Examples shared if the participant struggled: rule enforcement (e.g. anti harassment) and conflict resolution.
Interviewer: Why might it be necessary for a user to have full control over their chat platform data?
Examples shared if the participant struggled: data privacy, especially of sensitive information or personal conversations.
Interviewer: Based on thinking through these different situations, what level do you think is simultaneously the lowest possible and highest level of control necessary for chat platforms? Does this change in different contexts, e.g. personal or professional?